An electron moving along the x axis has a position given by x=16te-t m, where t is in seconds . How far is the electron from the origin when it momentarily stops ?

Grade:12th Pass

2 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
We can find the velocity of the electron being given by:
v=\frac{dx}{dt}=16\left ( e^{-t}-te^{-t} \right )
If the electron momentarily stops, then
v=0\Rightarrow 16e^{-t}\left ( 1-t \right )=0\Rightarrow t=1, \infty
Since t=\infty isnt possible, so we have t=1.
Thus, the distance from the origin would be given by:
x=16e^{-1} m
Regards
Sumit
Pushkala Krishnan
9 years ago
thank u . It didn’t strike me to differentiate .